# EEG-Driven 3D Object Reconstruction with Color Consistency and Diffusion Prior

### v1

@@ 5.2.3 Quantity
Among these metrics, 1), 2), and 3) are quantitative indicators for 2D images. As shown in Table 3, some values are missing due to the fact that certain metrics were not reported in the literature for the comparison methods, and the source code was not publicly available. For the Acc metric in the 50-way top-1 task, the accuracy of our proposed method is 21.4% higher than that of DreamDiffusion. In terms of the IS metric, our model also significantly outperforms Brain2Image and NeuroVision, indicating that the images generated by our model are more realistic and clearer. Compared to these models, our method demonstrates superior performance in both diversity and authenticity of the reconstructed images from EEG signals.
@@
Metrics 4) and 5) are qualitative indicators for 3D objects. As shown in Table 4, we evaluated the LPIPS and Contextual Distance metrics from different angles during the process of generating color-consistent 3D objects using EEG signals. The Acc metric here refers to the scores obtained using CLIP, where we assess the semantic similarity between the newly generated views and the reference views to measure the results.
@@
We conducted three ablation experiments to quantitatively analyze the impact of each component on the visual quality of color-consistent 3D objects. Table 5 presents the results of the ablation studies. When all components are utilized, our network achieves the best performance.
1) ${{\mathcal{L}}_{color}}$. To enable the EEG reconstruction of color-consistent 3D objects, we introduced neural style loss, which ensures that the 2D images reconstructed from different viewpoints using EEG maintain color consistency with GT. This approach facilitates the generation of color-consistent 3D objects.
2) ${{\mathcal{L}}_{EEG-Text}}$. Since the original Text-to-3D model aligns text prompts with novel views from different perspectives, we introduce the alignment between CLIP’s text encoder and EEG signals. ${{\mathcal{L}}_{EEG-Text}}$ aligns the text prompt of the rendered image $g(\theta)$ from a new viewpoint, the text prompt of the denoising image under diffusion priors, and the EEG signals.
3) ${{\mathcal{L}}_{region}}$. To enable the diffusion model with semantic region-aware capabilities, we keep the inherent ${{\mathcal{L}}_{ldm}}$ property of the LDM unchanged. By integrating latent EEG codes with regional images, we fine-tune the LDM, allowing it to reconstruct images with precise positional and spatial representation.
@@ 6 Conclusions
We propose an EEG-to-3D with Color Consistency approach, which, for the first time, utilizes EEG signals to generate high-fidelity 3D objects with consistent color. This method employs a two-stage strategy, integrating EEG signals, visual stimulus images, and style loss, to jointly fine-tune LDM and NeRF through collaborative training. The reconstructed 3D objects exhibit a high degree of realism and consistency in both geometry and color.
@@
At the Stage A phase, the semantic accuracy of 2D images reconstructed from EEG signals still has room for improvement, which is expected to further enhance the clarity of the generated 3D objects. Nevertheless, it is noteworthy that the findings of this study have demonstrated that brain signals can perceive and encode various visual features, such as the color, shape, and texture of objects.
Appendix A Main Generation Samples.
As shown in the figure below, we present several key reconstructed images, which feature color-consistent 3D objects. Our model utilizes EEG to directly reconstruct color-consistent 3D objects, accurately restoring the spatial positions of objects and the spatial relationships between multiple objects within the scene.
Prompt Visual Stimulus Images with ${{\mathcal{L}}_{color}}$ 3D Views (${20}^{o}$) ${45}^{o}$ ${70}^{o}$
Latent EEG Codes [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Latent EEG Codes [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Latent EEG Codes [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Prompt Visual Stimulus Images with ${{\mathcal{L}}_{color}}$ 3D Views (${20}^{o}$) ${45}^{o}$ ${70}^{o}$
Latent EEG Codes [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Latent EEG Codes [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Latent EEG Codes [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]

### v3

@@ 5.2.2 Quantity
Among these metrics, 1), 2), and 3) are quantitative indicators for 2D images. As shown in Table 3, some values are missing due to the fact that certain metrics were not reported in the literature for the comparison methods, and the source code was not publicly available. For the Acc metric in the 50-way top-1 task, the accuracy of our proposed method is 21.4% higher than that of DreamDiffusion  [13]. In terms of the IS metric, our model also significantly outperforms Brain2Image  [39] and NeuroVision  [40], indicating that the images generated by our model are more realistic and clearer. Compared to these models, our method demonstrates superior performance in both diversity and authenticity of the reconstructed images from EEG signals.
@@
Metrics 4) and 5) are qualitative indicators for 3D objects. As shown in Table 4, we evaluated the LPIPS and Contextual Distance metrics from different angles during the process of generating style-consistent 3D objects using EEG signals. The Acc metric here refers to the scores obtained using CLIP, where we assess the semantic similarity between the newly generated views and the reference views to measure the results.
@@
We conducted three ablation experiments to quantitatively analyze the impact of each component on the visual quality of style-consistent 3D objects. Table 5 presents the results of the ablation studies. When all components are utilized, our network achieves the best performance.
1) ${{\mathcal{L}}_{neuron-style}}$. To enable the EEG reconstruction of style-consistent 3D objects, we introduced neural style loss, which ensures that the 2D images reconstructed from different viewpoints using EEG maintain style consistency with GT. This approach facilitates the generation of style-consistent 3D objects.
2) ${{\mathcal{L}}_{EEG-CLIP}}$. Since the method proposed in this paper aligns EEG signals with novel views from different perspectives, we introduce the alignment between CLIP’s image encoder and EEG signals. ${{\mathcal{L}}_{EEG-CLIP}}$ aligns the novel view image $g(\theta)$ obtained under the diffusion prior with the EEG signals.
3) ${{\mathcal{L}}_{region}}$. To enable the diffusion model with semantic region-aware capabilities, we keep the inherent ${{\mathcal{L}}_{ldm}}$ property of the LDM unchanged. By integrating latent EEG codes with regional images, we fine-tune LDM, allowing it to reconstruct images with precise positional and spatial representation.
@@ 6 Conclusions
We propose an EEG-to-3D with style consistency approach, which, for the first time, utilizes EEG signals to generate high-fidelity 3D objects with consistent style. This method employs a two-stage strategy, integrating EEG signals, visual stimulus images, and style loss, to jointly fine-tune LDM and NeRF through collaborative training. The reconstructed 3D objects exhibit a high degree of realism and consistency in both geometry and color.
@@
At stage A phase, the semantic accuracy of 2D images reconstructed from EEG signals still has room for improvement, which is expected to further enhance the clarity of the generated 3D objects. Nevertheless, it is noteworthy that the findings of this study have demonstrated that brain signals can perceive and encode various visual features, such as the color, shape, and texture of objects.
As shown in Tables 6 and 7, we present several key reconstructed images featuring style-consistent 3D objects. Our model utilizes EEG to directly reconstruct these style-consistent 3D objects, accurately restoring the spatial positions of the objects and the spatial relationships between multiple objects within the scene. As illustrated in Table 8, the method proposed in this paper has been validated on the fMRI dataset, demonstrating the ability to reconstruct style-consistent 3D objects using fMRI data.
Visual Stimulus Images with ${{\mathcal{L}}_{neuron-style}}$ 3D Views (${20}^{o}$) ${45}^{o}$ ${70}^{o}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Visual Stimulus Images with ${{\mathcal{L}}_{neuron-style}}$ 3D Views (${20}^{o}$) ${45}^{o}$ ${70}^{o}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Visual Stimulus Images with ${{\mathcal{L}}_{neuron-style}}$ 3D Views (${20}^{o}$) ${45}^{o}$ ${70}^{o}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 8: Additional typical 3D objects generated by ours models based on the fmri dataset  [gao2023mind].
Figure 4: Our model reconstructs 3D objects using fMRI.